# Institutionalizing Folk Theories of Algorithms: How Multi-Channel Networks (MCNs) Govern Algorithmic Labor in Chinese Live-Streaming Industry

### v1

@@ 2.2. Algorithmic Management
Algorithmic management refers to the use of computational systems to allocate, supervise, and evaluate labor, increasingly replacing or supplementing traditional managerial functions in digital and platform-mediated work environments (Kellogg et al., 2020; McDaid et al., 2023; Möhlmann et al., 2021). Early scholarship in this field has focused on how platforms use algorithms to automate core managerial tasks, such as task assignment, performance monitoring, and discipline, especially in gig economy sectors like ride-hailing, delivery, and content moderation (Meijerink and Bondarouk, 2023; Momin, 2023). These studies highlight how algorithmic opacity and asymmetrical access to data reinforce structural inequalities, subjecting workers to rigid systems with limited transparency, accountability, or recourse (Rosenblat and Stark, 2016; Christin, 2020). Algorithms, in this view, become infrastructure for managerial control, embedding hierarchical logics in seemingly neutral technical systems.
Recent work has expanded the analytical lens beyond technical automation to include the social, affective, and discursive dynamics of algorithmic management. Scholars have documented how platforms govern not only through commands but through soft control, i.e., the subtle shaping of worker behavior via gamified dashboards, reward systems, performance metrics, and motivational narratives (Cameron, 2020; Jarrett, 2022; Wood and Lehdonvirta, 2023). These systems encourage workers to internalize platform values such as meritocracy, efficiency, and visibility, often prompting aspirational self-discipline under conditions of precarity. This turn in the literature emphasizes how algorithmic power operates at the level of subject formation, producing compliant, hopeful, and responsibilized workers even in the absence of overt coercion (Dasgupta et al., 2025; Cameron et al., 2021a).
However, much of this research has concentrated on the direct relationship between platforms and workers, with comparatively less attention to how intermediary institutions, such as third-party service providers, outsourcing firms, or managerial brokers, mediate, interpret, and operationalize algorithmic systems. These actors increasingly serve as epistemic and organizational bridges between opaque platforms and fragmented labor forces, playing a significant yet under-theorized role in shaping how algorithmic control is enacted on the ground. Our study contributes to this emerging line of inquiry by examining how such intermediaries not only interpret algorithmic logic but also strategically construct and institutionalize folk theories of algorithms to govern labor through belief, affect, and aspiration.
@@ 2.3. MCNs and Live-Streamers
This study investigates how folk theories of algorithms are strategically constructed, circulated, and institutionalized by intermediary organizations within China’s live-streaming industry. We focus on live-streaming industry as a particularly consequential site: streamers operate in highly volatile algorithmic environments, face extreme uncertainty in monetization, and often lack the technical knowledge to interpret the systems that shape their visibility and income (Sun et al., 2023; Yao et al., 2023; Hou et al., 2020; TikTok, 2022). These epistemic conditions make streamers especially dependent on externally provided explanations of how platform algorithms “work”, creating great context for the emergence and manipulation of folk theories.
@@
Unlike traditional gig work structured around discrete tasks, live-streamers must continuously perform for audiences in real time, often late into the night, while adapting their appearance, affect, and persona to meet platform and viewer expectations (Freeman and Wohn, 2020; Momin, 2023; Ye et al., 2023). The platform environment further exacerbates this instability. Streamers must rely on opaque algorithmic recommendation systems to gain visibility, yet they have little control or understanding of how these systems function (Sun et al., 2023; Yao et al., 2023). Success is contingent on being surfaced by algorithms, but algorithmic volatility means that visibility is uneven and unpredictable. Income models are primarily gift-based: most streamers do not receive a base salary, and instead depend on voluntary tips or fan gifts, the digital transactions that function more like micro-donations than stable compensation (Hou et al., 2020; Sun et al., 2023). With platforms retaining a significant portion of these gifts, and no obligation for viewers to contribute, streamers are left to shoulder both the risks and uncertainties of monetization (TikTok, 2022). Many live-streamers, particularly in the Global South or lower-income regions, enter this labor market with limited digital literacy and educational resources (Momin, 2023; Xiao et al., 2024a; Tang et al., 2022). They often lack the technical background to interpret algorithmic behavior, analyze platform metrics, or optimize engagement strategies. This limited algorithmic understanding further amplifies their precarity, making them more vulnerable to misinformation, overwork, and self-blame when platform performance falters (Uttarapong et al., 2021; Ye et al., 2023).
@@
In this context, MCNs have become central to the construction of algorithmic knowledge and the mediation of labor, which has been a great case for our study in institutionalized folk theories of algorithms. Originating in Western markets such as YouTube, the MCN model initially offered monetization support, branding strategy, and analytics services to individual creators and channel collectives (Lobato, 2016; Postigo, 2016). While the influence of MCNs on YouTube has fluctuated, research has shown that these entities helped consolidate and industrialize early creator labor by aggregating talent, mediating intellectual property claims, and standardizing aesthetic and technical norms across channels (Cunningham and Craig, 2019; Chen et al., 2024). Even when not formally embedded within platform governance, MCNs in the West have played significant roles in disciplining creative labor through soft control mechanisms such as branding scripts, content quotas, and algorithm-focused consultations (Siciliano, 2023).
@@
We selected China as our main study context. In China, MCNs have become far more structurally embedded within the operational logics of platforms like Douyin, Kuaishou, and Huya. Rather than serving merely as optional support services, MCNs in China have been institutionalized into platform governance frameworks, actively assisting platforms in managing, monitoring, and optimizing the labor of live-streamers (Zhang and Tong, 2024; Liang and Ji, 2024; Liu et al., 2023). Unlike YouTube and other Western live-streaming platforms, where most creators remain relatively autonomous  (Alexander, 2018), Chinese live-streaming platforms rely heavily on MCNs to coordinate recruitment, training, and day-to-day performance evaluation. As a result, independent live-streamers are increasingly rare, and platform labor is mediated through MCNs that function as both infrastructural supports and informal managers (Gardner and Lehnert, 2016). These MCNs often promote their technical and analytic capacity to “manufacture” success by decoding and exploiting algorithmic mechanisms, branding themselves as algorithm-savvy “talent incubators” (Zhang and Tong, 2024; Zhang et al., 2024).
@@
MCNs act as interpretive infrastructures in the algorithmic era. Their authority stems not from employment arrangements but from their ability to decode opaque platform systems, train streamers to meet algorithmic expectations, and define what counts as “desirable” labor. This role is especially critical in China’s live-streaming industry, where many streamers, often rural migrants or young workers with limited digital literacy (Zhao, 2024; Tang et al., 2022; Xiao et al., 2024a), rely on MCNs as their main, sometimes only, source of guidance. Zhang and Tong recently offered a systematic account of the historical development of the MCN ecosystem in China and argued that Chinese content platforms highly rely on MCNs to absorb the risks of professional content production  (Zhang and Tong, 2024). They also acknowledge that MCNs play an active role in helping creators become familiar with platform algorithms and optimize their content accordingly  (Zhang and Tong, 2024). However, they stop short of investigating a critical question: how do MCNs specifically construct and disseminate knowledge about algorithms, and why do they do so in particular ways? As a result, the role of MCNs in shaping creators’ epistemic relationship to platform algorithms as both interpreters and enforcers remains undertheorized. Our study addresses this gap by exploring how MCNs strategically mobilize informal algorithmic knowledge not only to support creators, but also to discipline and manage them.
@@ 3. Methods
In this study, we conducted extensive fieldwork primarily in Beijing and Changsha in China, focusing on MCN companies and live-streamers. From April to December 2023, we engaged in a nine-month in-depth ethnographic investigation. Our fieldwork involved observing the daily operations of three MCN companies (AB, QC, XH in our cases)—pseudonyms used to protect the anonymity of participating organizations—including how they set up live-streaming environments, provided technical and emotional support, and managed interactions among managers, technologists, and live-streamers. We also followed streamers’ broader daily routines to understand their labor conditions and platform practices in context. All observations were conducted with the consent and cooperation of both the MCN companies and the individual participants.
@@
To supplement this ethnography, we conducted one-on-one online interviews with 24 live-streamers and 13 MCN practitioners, each lasting approximately one hour. While our immersive fieldwork was based in three MCNs, our interview sample expanded to include participants from a total of seven MCN companies (AB, QC, XH, KX, HG, RZ, XY).
@@
This research received approval from the university’s Institutional Review Board (IRB). Throughout the study, we were transparent about our researcher roles and obtained informed consent from all participants. We also engaged in ongoing reflexive practices, including regular internal discussions among research team members to critically examine power dynamics, positionality, and potential sources of bias.
3.1. Participants: Live-Streamers (N=24), MCNs Practitioners (N=13)
@@
We interviewed 24 live-streamers (L1–L24) active on major Chinese platforms including Laifeng, MoMo, YY, Douyin, and Huajiao. Participants ranged in age from 21 to 34 and came from diverse gender, educational, and geographic backgrounds (see Table 1). About two-thirds (n=15) were full-time streamers, while the remainder worked part-time due to unstable income. We also interviewed 13 MCN practitioners (M1–M13) from seven companies spanning different stages of organizational development, including startups, growth-stage firms, and mature MCNs (see Table 2). Their roles included executives (n=5), managers (n=6), and technicists (n=3), with one participant holding a hybrid manager-technician position. Executives made strategic decisions; managers coordinated scheduling and performance; and technicists provided technical support and training. We intentionally sampled  (Campbell et al., 2020) participants across a range of employment statuses, platform types, MCN development stages, and institutional roles to reflect the common experience of China’s live-streaming industry. By including voices from both workers and MCNs, our study design aims to surface not only individual experiences but also broader structural patterns of visibility, precarity, and control.
@@
Name Gender Age MCN Work Tenure Education Platform Fans Employment Monthly Income
L1 Female 21 AB 8 months Undergraduate Laifeng 3k Part-time ¥1000-2000 / \$140-280
L2 Male 24 AB 6 months Master Laifeng 2k Part-time ¥1000-1500 / \$140-210
L3 Female 21 KX 3 months Undergraduate Laifeng 1k Part-time ¥3000-4000 / \$420-560
L4 Female 25 QC 3 months Master Laifeng 1k Part-time ¥2000-3000 / \$420-700
L5 Female 22 XH 7 months Senior high school Laifeng 23k Full-time ¥2000-4000 / \$280-560
L6 Female 26 XH 8 months Senior high school Laifeng 1k Full-time ¥1500-2500 / \$210-350
L7 Male 24 QC 3 months Junior college Laifeng 1k Full-time ¥1500-3000 / \$210-420
L8 Female 24 QC 9 months Junior college Laifeng 1k Part-time ¥2000 / \$280
L9 Female 18 KX 2 years Junior college Laifeng 6k Full-time ¥3500-4000 / \$490-560
L10 Female 23 XH 5 years Senior high school Laifeng 148k Full-time ¥10000-15000 / \$1400-2100
L11 Female 26 XH 4 years Undergraduate Laifeng 98k Full-time ¥10000-12000 / \$1400-1680
L12 Female 28 XH 2 years Senior high school Laifeng 57k Full-time ¥8000-10000 / \$1120-1400
L13 Female 20 XY 3 years Junior college MoMo 12k Full-time ¥3000-6000 / \$420-840
L14 Female 28 XY 1 year Senior high school MoMo 1k Full-time ¥4000-4500 / \$560-630
L15 Female 26 XY 1.5 years Junior college MoMo 1k Full-time ¥3000-5000 / \$420-700
L16 Female 24 RZ 2.5 years Junior college $Y$ 2k Full-time ¥1000-3500 / \$140-490
L17 Female 28 RZ 1 year Undergraduate $Y$ 1k Full-time ¥2000-4000 / \$240-560
L18 Female 23 XH 1 year Undergraduate Douyin 2k Full-time ¥1500-2000 / \$210-280
L19 Male 21 XH 1.5 years Undergraduate Douyin 83k Part-time ¥2000-3000 / \$280-420
L20 Male 24 RZ 2 years Undergraduate Douyin 75k Part-time ¥3000-4000 / \$420-560
L21 Male 34 HG 3 months Undergraduate Douyin 2k Part-time ¥1000-1500 / \$140-210
L22 Male 26 HG 4 months Master Douyin 9k Part-time ¥2000-2500 / \$280-350
L23 Male 29 RZ 2 years Junior college Douyin 15k Full-time ¥4000-5000 / \$560-700
L24 Female 32 RZ 1 year Master Huajiao 3k Full-time ¥3000-3500 / \$420-490
@@
M1 QC Mature Executive Male 42 4 years Master
M2 KX Startup Manager Female 35 2 years Junior college
M3 AB Mature Executive Female 42 5 years Undergraduate
M4 XH Growth Manager/Technicist Male 40 4 years Junior college
M5 AB Mature Manager Female 40 1 year Undergraduate
M6 XH Growth Executive Male 48 5 years Junior college
M7 RZ Startup Executive Female 46 2 years Ph.D
M8 HG Startup Manager Male 53 2 years Undergraduate
M9 XY Mature Manager Male 25 2 years Junior college
M10 RZ Startup Manager Female 35 2 years Master
M11 XY Mature Executive Male 30 3 years Junior college
M12 XH Growth Technicist Male 24 1 year Junior college
M13 XH Growth Technicist Male 22 6 months Undergraduate

### v2

@@ 2.2. Algorithmic Management
Algorithmic management refers to the use of computational systems to allocate, supervise, and evaluate labor, increasingly replacing or supplementing traditional managerial functions in digital and platform-mediated work environments (Kellogg et al., 2020; McDaid et al., 2023; Möhlmann et al., 2021). Early scholarship in this field has focused on how platforms use algorithms to automate essential managerial procedures, such as task assignment, performance evaluation, and behavior discipline, especially in gig economy sectors like ride-hailing, delivery, and content moderation (Meijerink and Bondarouk, 2023; Momin, 2023). These studies highlight how algorithmic opacity and asymmetrical access to data reinforce structural inequalities, subjecting workers to rigid systems with limited transparency (Rosenblat and Stark, 2016; Christin, 2020). Algorithms, in this view, become infrastructure for managerial control, embedding hierarchical logics in seemingly neutral technical systems.
Recent work has expanded the analytical lens beyond technical automation to include the social, affective, and discursive dynamics of algorithmic management. Scholars have documented how platforms govern not only through direct spoken commands but soft control, i.e., the subtle shaping of worker behavior via gamified dashboards, reward systems, performance metrics, and motivational narratives (Cameron, 2020; Jarrett, 2022; Wood and Lehdonvirta, 2023). These systems encourage workers to internalize platform values such as meritocracy, efficiency, and visibility, often prompting aspirational self-discipline under conditions of precarity. This turn in the literature emphasizes how algorithmic power operates at the level of subject formation, producing compliant, hopeful, and responsibilized workers even in the absence of overt coercion (Dasgupta et al., 2025; Cameron et al., 2021a).
However, much of this research has concentrated on the direct relationship between platforms and workers, with comparatively less attention to how intermediary institutions, such as third-party service providers, outsourcing firms, or managerial brokers mediate algorithmic systems. These actors have been increasingly serving as epistemic and organizational bridges between opaque platforms and fragmented labor forces, playing a significant yet under-theorized role in shaping how algorithmic control is enacted on the ground.
@@ 2.3. MCNs and Live-Streamers
This study investigates how interpretations of algorithms are strategically constructed by intermediary organizations within China’s live-streaming industry. We focus on live-streaming industry as a particularly consequential site: streamers operate in highly volatile algorithmic environments, face extreme uncertainty in monetization, and often lack the technical knowledge to interpret the systems that shape their visibility and income (author = John Peikang Sun et al., 2023; Yao et al., 2023; Hou et al., 2020; TikTok, 2022). These epistemic conditions make streamers especially dependent on externally provided explanations of how platform algorithms “work”, creating context for the emergence and manipulation of algorithmic interpretations.
@@
Unlike traditional gig work structured around discrete tasks, live-streamers have to continuously perform for audiences in real time, often late into the night, while adapting their appearance, affect, and persona to meet platform and audience expectations (Freeman and Wohn, 2020; Momin, 2023; Ye et al., 2023). The platform environment further exacerbates this instability. Streamers have to rely on opaque algorithmic recommendation systems to gain visibility, yet they have little security, knowledge or control of how these systems function (author = John Peikang Sun et al., 2023; Yao et al., 2023). Success is contingent on being surfaced by algorithms, but algorithmic volatility means that visibility is uneven and unpredictable. Income mechanisms are primarily gift-based: most streamers do not receive a base salary, and instead depend on voluntary tips or fan gifts from the audience, meaning the digital transactions function more like micro-donations than stable compensation (Hou et al., 2020; author = John Peikang Sun et al., 2023). With platforms retaining a significant portion of these gifts, and no obligation for viewers to contribute, streamers are left to bear the risks and uncertainties of monetization (TikTok, 2022). Many live-streamers, particularly in the Global South or lower-income regions, enter this labor market with limited digital literacy and educational resources (Momin, 2023; Xiao et al., 2024a; Tang et al., 2022). They often lack the technical background to interpret algorithmic behavior, analyze platform metrics, or optimize engagement strategies. This limited algorithmic understanding further amplifies their precarity, making streamers vulnerable to misinformation, overwork, self-blame, and self-discipline when platform performance falters (Uttarapong et al., 2021; Ye et al., 2023).
@@
In this context, MCNs have become central to the construction of algorithmic knowledge and the mediation of labor, which has been a great case for our study in the discursive construction of algorithms. Originating in Western markets such as YouTube, the MCN model initially offered monetization support, branding strategy, and analytics services to individual creators and channel collectives (Lobato, 2016; Postigo, 2016). While the influence of MCNs on YouTube has fluctuated, research has shown that these entities helped consolidate and industrialize early creator labor by aggregating talent, mediating intellectual property claims, and standardizing aesthetic and technical norms across channels (Cunningham and Craig, 2019; Chen et al., 2024). Even when not formally embedded within platform governance, MCNs in the West have played significant roles in disciplining creative labor through soft control mechanisms such as branding scripts, content quotas, and algorithm-focused consultations (Siciliano, 2023).
@@
We selected China as our main context to focus based on it’s rapidly developing live-streaming industries and burgeoning intermediary companies. In China, MCNs have become more structurally embedded within the operational logic of platforms such as Douyin, Kuaishou, and Huya. In the meantime, MCNs in China have been institutionalized into platform governance frameworks, actively assisting platforms in managing, monitoring, and optimizing the production of live-streamers (Zhang and Tong, 2024; Liang and Ji, 2024; Liu et al., 2023; Wang and Cao, 2024). Unlike YouTube and other western live-streaming platforms, where most creators remain relatively autonomous  (Alexander, 2018), Chinese live-streaming platforms rely heavily on MCNs to coordinate recruitment, training, and day-to-day performance evaluation. As a result, independent live-streamers experience higher competition, and platform labor mediated through MCNs functions both as infrastructural supports and informal managers (Gardner and Lehnert, 2016). These MCNs often promote their technical and analytic capacity to “manufacture” success by decoding and exploiting algorithmic mechanisms, branding themselves as algorithm-savvy “talent incubators” (Zhang and Tong, 2024; Zhang et al., 2024).
@@
MCNs act as interpretive infrastructures in the algorithmic era. Their authority stems not from employment arrangements but from their ability to decode opaque platform systems, train streamers to meet algorithmic expectations, and define what counts as desirable labor. Zhang and Tong recently offered a systematic account of the historical development of the MCN ecosystem in China and argued that Chinese content platforms highly rely on MCNs to absorb the risks of professional content production  (Zhang and Tong, 2024). They also acknowledge that MCNs play an active role in helping creators become familiar with platform algorithms and optimize their content accordingly  (Zhang and Tong, 2024). However, there is still a gap in literature on investigating critical questions: how do MCNs specifically construct and disseminate knowledge about algorithms, and why do they do so in particular ways? Thus, our research aims to explore the role of MCNs in shaping creators’ epistemic relationship to platform algorithms as both interpreters and enforcers.
@@ 3. Methods
In this study, we conducted extensive fieldwork primarily in Beijing and Changsha in China, focusing on MCN companies and live-streamers. From April to December, 2023, we engaged in a nine-month in-depth ethnographic investigation. Our fieldwork involved observing the daily operations of three MCN companies (A, B, C in our cases)—pseudonyms used to protect the anonymity of participating organizations—including how they set up live-streaming environments, provided technical and emotional support, and managed interactions among managers, technologists, and live-streamers. We also followed streamers’ broader daily routines to understand their labor conditions and platform practices in context. All observations were conducted with the consent and cooperation of both the MCN companies and the individual participants.
@@
To supplement this ethnography, we conducted one-on-one online interviews with 24 live-streamers, seven former live-streamers, and 13 MCN practitioners. While our immersive fieldwork was based in three MCNs, our interview sample expanded to include participants from a total of seven MCN companies (A, B, C, D, E, F, G).
@@
This research received approval from the university’s Institutional Review Board (IRB). Throughout the study, we were transparent about our researcher roles and obtained informed consent from all participants.
3.1. Participants: Live-Streamers (N=24), Former Live-Streamers (N=7), MCNs Practitioners (N=13)
@@
We interviewed 24 live-streamers (L1–L24) active on major Chinese platforms including Laifeng, MoMo, YY, Douyin, and Huajiao. Participants ranged in age from 21 to 34 and came from diverse gender, educational, and geographic backgrounds (see Table 1). About two-thirds (n = 15) were full-time streamers, while the remainder worked part-time, supplementing their unstable streaming income with jobs in nail salons, barbershops, massage parlors, supermarkets, electronics factories, or as fitness instructors, models, and salespeople. We also interviewed 13 MCN practitioners (M1–M13) from seven companies spanning different stages of organizational development, including startups, growth-stage firms, and mature MCNs (see Table 2). Their roles included executives (n=5), managers (n=6), and technicists (n=3), with one participant holding a hybrid manager-technician position. Executives made strategic decisions; managers coordinated scheduling and performance; and technicists provided technical support and training. We also interviewed seven former live-streamers (F1–F7) who had previously been affiliated with MCNs and had left the live-streaming industry (see Table 3). It is noted that F1 and F7 initially left their MCNs to stream independently without any MCNs for a while before ultimately exiting the industry. F5 had switched between two MCN companies prior to departure.
As of December 2023, approximately 15.08 million individuals in China have adopted live-streaming as their primary occupation, with 57.1% citing economic income as their main motivation for entering the industry. 68% of professional live-streamers report average monthly incomes below 3,000 RMB (Approximately \$420) (People’s Daily Online, [n. d.]; China Performance Industry Association, Online Performance Branch and Kuaishou([n. d.]), Live Streaming).
@@
Name Gender Age MCN Work Tenure Education Platform Fans Employment Income Domain
L1 Woman 21 A 8 months Undergraduate Laifeng 3k Part-time ¥1000-2000 Chat
L2 Man 24 A 6 months Master Laifeng 2k Part-time ¥1000-1500 Chat
L3 Woman 21 D 3 months Undergraduate Laifeng 1k Part-time ¥3000-4000 Music
L4 Woman 25 B 3 months Master Laifeng 1k Part-time ¥2000-3000 Music
L5 Woman 22 C 7 months Senior high school Laifeng 23k Full-time ¥2000-4000 Music
L6 Woman 26 C 8 months Senior high school Laifeng 1k Full-time ¥1500-2500 Dance
L7 Man 24 B 3 months Junior college Laifeng 1k Full-time ¥1500-3000 Chat
L8 Woman 24 B 9 months Junior college Laifeng 1k Part-time ¥2000 Chat
L9 Woman 18 D 2 years Junior college Laifeng 6k Full-time ¥3500-4000 Chat
L10 Woman 23 C 5 years Senior high school Laifeng 148k Full-time ¥10000-15000 Music
L11 Woman 26 C 4 years Undergraduate Laifeng 98k Full-time ¥10000-12000 Music
L12 Woman 28 C 2 years Senior high school Laifeng 57k Full-time ¥8000-10000 Dance
L13 Woman 20 G 3 years Junior college MoMo 12k Full-time ¥3000-6000 Chat
L14 Woman 28 G 1 year Senior high school MoMo 1k Full-time ¥4000-4500 Chat
L15 Woman 26 G 1.5 years Junior college MoMo 1k Full-time ¥3000-5000 Chat
L16 Woman 24 F 2.5 years Junior college $Y$ 2k Full-time ¥1000-3500 Music
L17 Woman 28 F 1 year Undergraduate $Y$ 1k Full-time ¥2000-4000 Music
L18 Woman 23 C 1 year Undergraduate Douyin 2k Full-time ¥1500-2000 Music
L19 Man 21 C 1.5 years Undergraduate Douyin 83k Part-time ¥2000-3000 Music
L20 Man 24 F 2 years Undergraduate Douyin 75k Part-time ¥3000-4000 Fitness
L21 Man 34 E 3 months Undergraduate Douyin 2k Part-time ¥1000-1500 Music
L22 Man 26 E 4 months Master Douyin 9k Part-time ¥2000-2500 Chat
L23 Man 29 F 2 years Junior college Douyin 15k Full-time ¥4000-5000 Dance
L24 Woman 32 F 1 year Master Huajiao 3k Full-time ¥3000-3500 Music
@@
M1 B Mature Executive Man 42 4 years Master
M2 D Startup Manager Woman 35 2 years Junior college
M3 A Mature Executive Woman 42 5 years Undergraduate
M4 C Growth Manager/Technicist Man 40 4 years Junior college
M5 A Mature Manager Woman 40 1 year Undergraduate
M6 C Growth Executive Man 48 5 years Junior college
M7 F Startup Executive Woman 46 2 years Ph.D
M8 E Startup Manager Man 53 2 years Undergraduate
M9 G Mature Manager Man 25 2 years Junior college
M10 F Startup Manager Woman 35 2 years Master
M11 G Mature Executive Man 30 3 years Junior college
M12 C Growth Technicist Man 24 1 year Junior college
M13 C Growth Technicist Man 22 6 months Undergraduate
Table 3. Information of Former Live-streamers Participants.
Name Gender Age MCN Work Tenure Education Platform Fans Employment Income Domain
F1 Woman 22 / 1 year Undergraduate Douyin 34k Full-time ¥4000-5000 Chat
F2 Woman 31 A 6 months Undergraduate Laifeng 2k Part-time ¥2000-3000 Chat
F3 Woman 28 F 1.5 years Senior high school $Y$ 6k Full-time ¥3000-5000 Music
F4 Woman 26 F 8 months Senior high school $Y$ 5k Full-time ¥3000-3500 Dance
F5 Man 29 G 2 years Senior high school Momo 23k Full-time ¥4000-5000 Chat
F6 Man 23 E 6 months Junior college Douyin 52k Part-time ¥1000-3000 Fitness
F7 Man 21 / 6 months Undergraduate Douyin 28k Part-time ¥2000-3000 Fitness
Note. Live-streaming data reflects account information at the time of departure from the industry.